OpenStudy (anonymous):

Given the function f(x) = 3x + 1, evaluate f(a + 1). A. 3a + 1 B. 3a + 2 C. 3a + 4

OpenStudy (matt101):

f(a+1) = 3(a + 1) + 1 = 3a + 4

OpenStudy (callisto):

f(a + 1) = 3(a + 1) + 1= 3a+4 =C